OpenStudy (anonymous):

Fahrenheit and Celsius temperatures are related by the formula f = 9/5 c + 32 What is the Celsius temperature if the temperature is -22ºF? a. -30ºC c. 5ºC b. -7.6ºC d. 19ºC

OpenStudy (anonymous):

F=(9/5)C +32 F-32 =(9/5) C (F-32)(5/9)=C

OpenStudy (anonymous):

now, F=-22 C= (-22-32)(5/9) =(-54)(5/9)= -30 thus the answer is option a

OpenStudy (anonymous):

thank you